OpenStudy (anonymous):

what is a polynomial function in standard form with zeros 1, 2, -3, and -1? g(x)= x^4 + x^3 - 7x^2-x+6 g(x)= x^4 + x^3 + 7x^2 - x + 6 g(x)= x^4 + x^3 - 7x^2 - x - 6 g(x)= x^4 - x^3 - 7x^2 - x + 6

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@iambatman

OpenStudy (mathmate):

A polynomial with zeroes a,b,c...z looks typically like: f(x)=(x-a)(x-b)(x-c)....(x-z). You could put an optional multiplicative constant but that wont change the zeroes. Example: f(x)=(x-2)(x-4) has roots x=2 and x=4. Multiply them out gives f(x)=x^2-6x+8